OpenStudy (anonymous):

5x^3 + 6x + 8 by x + 2 a. 5x -10 + 26/(x + 2) b. -44 c. 5x2 - 10x + 26 d. 5x2 - 10x + 26 -44/(x + 2)

OpenStudy (anonymous):

divided by?

OpenStudy (anonymous):

divide by x + 2

OpenStudy (anonymous):

and you want the remainder i guess

OpenStudy (anonymous):

it is 44

OpenStudy (anonymous):

your divisor is gonna be -2 for this one

OpenStudy (anonymous):

actually it is -44

OpenStudy (anonymous):

but if you do not want the remainder then i am not sure what the question is asking

OpenStudy (anonymous):

d.

jhonyy9 (jhonyy9):

5x*3 +6x +8 / x+2 = 5x*2 -10x +26 5x*3 +10x*2 ______________ ---- -10x*2 +6x -10x*2 -20x ------------ ------ 26x +8 26x +52 --------- ---- -44

OpenStudy (anonymous):

7/2 = 3+1/2 From Mathematica Home Edition: PolynomialQuotientRemainder[5 x^3 + 6 x + 8, x + 2, x] = {26 - 10 x + 5 x^2, -44} or\[5 x^2-10 x+26+\frac{-44}{x+2}\to 5 x^2-10 x+26-\frac{44}{x+2} \]